OpenStudy (anonymous):

what is the solution to the system of equations |{ x-2y+z=10 <|{ z=6 |{ 3x-2y+2z=8 the possible answers are (0,-2,6) (0,2,6) (-1,2,6) (0,2,-6)

OpenStudy (amistre64):

none of those answers make any sense with the stated problem. You can dbl check them by working the backwards; plug the values into the equations and see if they are true.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

its the first one?

OpenStudy (amistre64):

x-2y+z=10 <-- (0,2,6) makes this false z=6 <-- (0,2,-6) makes this false 3x-2y+2z=8 <-- (0,-2,6) makes this false

OpenStudy (amistre64):

x-2y+z=10 ; (-1,2,6)? -1 -2(2) +6 = 10 -1 -4 +6 = 10 -5 +6 = 10 .. FALSE

OpenStudy (amistre64):

none of them satisfy all the equations

OpenStudy (anonymous):

but the first one satisfies two out of three so thats the best answer

OpenStudy (anonymous):

it's 0,2,6!!!!!